14. The present age of a father is equal to the sum of the ages of his 5 children. 12 years hence, the sum the ages of his children will be twice the ages of their father. Find the present age of the father.
Answer by ankor@dixie-net.com(15661)

The present age of a father is equal to the sum of the ages of his 5 children.
12 years hence, the sum the ages of his children will be twice the ages of their father.
Find the present age of the father.
:
let f = father's present age
let s = sum of the 5 children's now
then
f = s
"12 years hence, the sum the ages of his children will be twice the ages of their father."
(12 yrs hence, add 12 to each of the five children)
2(f + 12) = s + 5(12)
2f + 24 = s + 60)
2f = s + 60 - 24
2f = s + 36
replace s with f
2f = f + 36
2f - f = 36
f = 36 yrs is the father's present age